### v1

@@ 4.1 Experimental Setup
In video glance process, we uniformly sample 4 frames. If viewing the entire video is required, we uniformly sample 32 frames for processing. In the Feature Extraction stage, we sample the video at 1 frame per second (1 FPS) for feature extraction. We use pretrained CLIP-ViT-B/32 [36] as the feature extractor, and the extracted features are used for clustering. In Shot Slection stage, we sample 16 frames from each shot as input and use a fine-tuned LongCLIP-B [56] to calculate shot-text similarity. The fine-tuning details and hyperparameters are described in the supplementary doc. We retrieve shots with similarity scores above the 0.8 threshold. In the 1st round, we select top 1 shot and sample 16 frames as initial frames at 1st round. In each subsequent round, we select top 2 shots, and divide each shot into 2 subshots. Then, we sample 8 frames from each subshot, and add them into frame set. In the Shot Partition stage, the number of clusters $K$ is set to 6 in 1st round and 2 in subsequent rounds. In the Shot Reflection stage, If the confidence score is greater than 2, the answer is considered correct; otherwise, the next iteration round continues until the confidence requirement is met or the maximum number of iterations is reached. The maximum number of iterations is set to 3. All implementations are based on Pytorch with 2 A800-80G GPUs.
4.2 Comparison with SOTA
Performance Analysis.
As shown in Table 1, we evaluate VideoChat-A1 on four mainstream benchmarks. Specifically, VideoChat-A1 significantly outperforms open-source 7-8B models, comparable to or exceeding that of large open-source or closed-source models. For example, on EgoSchema, VideoChat-A1 (InternVL2.5-8B) achieves 72.1%, surpassing InternVideo2.5-8B and VideoTree by 8.2% and 11.0%. On LongVideoBench, VideoChat-A1 (InternVideo2.5-8B) achieves 65.4%, surpassing ShareGPT4Video-8B and Gemini 1.5Pro by 25.7% and 1.4%. On MLVU, VideoChat-A1 (InternVideo2.5-8B) achieves 76.2%, outperforming GPT-4o by 11.6%. On VideoMME, VideoChat-A1 (InternVL2.5-8B) achieves 72.3% (w/o subs) and 77.0 (w/ subs), outperforming InternVL-2.5-78B by 0.2% and 3.0%.
Efficiency Analysis.
@@ 4.2 Comparison with SOTA
Table 2 presents a comparison of the average frame number and average inference time across various models. Compared to both closed-source models and large open-source models, VideoChat-A1 achieves substantial gains in efficiency. For instance, VideoChat-A1 (InternVL2.5-8B) processes only 35.9 frames and completes inference in 14.7 seconds on average, reducing frame usage by over 90% and inference time by more than 85% compared to GPT-4o, Gemini 1.5 Pro, and Qwen2.5-VL-72B. Despite these significant savings, our method maintains competitive performance across benchmarks. In contrast to small open-source models such as Qwen2.5-VL-7B, VideoChat-A1 not only delivers significantly better performance but also requires fewer video frames and achieves faster or comparable inference.
4.3 Ablation Study
@@
Table 3: Ablation on the effect of Video Glance in VideoChat-A1.
@@
In this study, we conducted an ablation study to verify the impact of the video glance on the performance of the model. Table 3 shows that the model with the video glance achieves higher scores on all the tested tasks. The improvements proving the effectiveness of video glance.
@@
As shown in Table 5, we conducted an ablation study on VideoChat-A1 (Qwen2.5-VL-7B) to evaluate the impact of Shot Partition (SP) and Chain-of-Shot (CoS) across Egoschema, LongVideoBench, MLVU, and VideoMME datasets. The Egoschema and LongVideoBench are abbreviated as Ego and LVbench in the table because of the limited space. For a fair comparison, in the setup without shot partition, we evenly divided the entire video into six shots on average in the 1st round. In subsequent rounds, for the selected shots that need to be expanded, they are directly partitioned into two shots on average. Regarding the setup without Chain-of-Shot, we removed the Chain-of-Shot mechanism. Instead, the model was made to answer three rounds. In each round, the model samples additional frames from the video and incorporates them into the existing frame set from the previous round. Then we conducted a vote, selecting the answer that appeared most frequently as the final response. Individually, enabling shot partition or Chain-of-Shot boosts performance. Their combined use yields the best results. For example, on MLVU, the score rises from 66.2 (with neither) to 71.9 (both enabled), highlighting their complementary roles in enhancing long-video understanding.
SP
CoS Ego LVBench MLVU VideoMME
@@
Table 4: Ablation of Shot Partition(SP) and Chain-of-Shot(CoS) in VideoChat-A1.
@@
Table 5: Ablation of different method of shot partition in VideoChat-A1 (Qwen2.5-VL-7B).
@@
Table 7: Ablation over the number of initialized shots in VideoChat-A1 (Qwen2.5-VL-7B).
@@
As shown in Table 5, we evaluate the impact of three methods of shot partition across multiple benchmarks. Cluster Partition refers to the clustering approach used in VideoTree [46], where all frames within each cluster are treated as a single shot, even though they may not form a continuous video segment. Average Partition refers to dividing each shot into smaller shots by uniformly splitting it into equal parts. Our proposed Shot Partition method consistently outperforms both Cluster Partition and Average Partition across all four benchmarks. These results demonstrate that partitioning videos into temporally coherent shots based on semantic information leads to significantly better understanding of long videos. Our Shot Partition achieves best accuracies of 70.7, 64.2, 71.9, and 71.8/76.5 on EgoSchema, LongVideoBench, MLVU, and VideoMME. Compared to the better of the other two methods, our Shot Partition method achieves gains of 1.6%, 5.8%, 4.2%, and 4.0%/3.7%
@@
We evaluate the impact of varying the maximum number of reflection rounds across multiple benchmarks. As shown in Table 7, increasing the number of rounds consistently leads to performance improvements on all the tested dataset. For example, the accuracy on EgoSchema increases from 67.3% when using one reflection round to 70.7% with three rounds, while the accuracy on LongVideoBench improves from 60.1% to 64.2%. In the context of multi-round discussions, increasing the number of discussion rounds tends to enhance the effectiveness. After three rounds of the chain-of-shot process, the results are already quite satisfactory. As the fourth round yields only marginal improvements, we ultimately adopt the three-round setting to balance performance gains and computational cost.
@@
To evaluate the impact of the number of initial shots on the experimental results, the initial shot count is set to 3, 6, and 9, respectively. As shown in Table 7, the 6-shot configuration achieves best accuracies of 70.7, 64.2, 71.9, and 71.8/76.5 on EgoSchema, LongVideoBench, MLVU, and VideoMME, respectively. These correspond to gains of 1.6%, 5.8%, 4.2%, and 4.0%/3.7% over the better of the other two configurations, demonstrating that the 6-shot setting yields the best overall performance.
5 Conclusion
@@
In this paper, we introduced VideoChat-A1, a novel agent-based framework for effectively addressing long video understanding tasks. Unlike existing approaches that often overlook the intrinsic shot-based structure of long videos, our method explicitly employs a chain-of-shot reasoning paradigm. Specifically, VideoChat-A1 progressively identifies and refines relevant video segments through three key iterative steps: Shot Selection, Shot Partition, and Shot Reflection. This iterative interaction mechanism closely mimics human cognitive processes, allowing the agent to thoughtfully reason and accurately answer user questions from long videos. Extensive experimental results demonstrate the effectiveness and efficiency of our approach, highlighting the substantial benefits of our method in long video understanding tasks.
Appendix A Prompts of VideoChat-A1.
In this section, we introduce the prompts used in VideoChat-A1, as shown in Table LABEL:tab:prompts. These prompts are carefully designed to precisely fulfill the requirements of each step in the VideoChat-A1 pipeline.
Table 8: The prompting templates used in different steps of VideoChat-A1.
Prompt Content
Prompt to decide whether to watch the whole video You are given a single-choice question, options, subtitles, and some frames of the long video. You should not only look at the textual information but also consider the input visual information, taking everything into account. If you can answer the question accurately and comprehensively based on the existing information, especially the visual information, and further watching the entire video will not significantly improve the quality of the answer, then you don’t need to watch the entire video and can answer ’No.’. However, if the existing information is not sufficient to fully answer the question, and watching the entire video may obtain information crucial for answering the question, please reply ’Yes’.
The frame tokens: {Frame tokens}
The question is:{Question}
The options is:{Options}
The subtitiles is:{Subtitiles}(If have)
Output:{Yes/No}
Prompt for generate the key information Given some frames from a long video, subtitles, a single-choice question, and options, identify the key information needed to answer the question. Focus on visual cues, context, and temporal relationships within the frames. Limit your response to 50 words.
The frame tokens: {Frame tokens}
The question is:{Question}
The options is:{Options}
The subtitiles is:{Subtitiles}(If have)
Prompt for generating the answer Given some frames from a long video, subtitles, a single-choice question, and options, select the best answer to the following question based on the video and the subtitles. Respond with only the letter (A, B, C, or D) of the correct option.
The frame tokens: {Frame tokens}
The question is:{Question}
The options is:{Options}
The subtitiles is:{Subtitiles}(If have)
The Key Info is:{Key Info}
The best answer is:
Prompt for generating the reason You are given a single-choice question, options, some frames of the long video, the key information of the question and the choice you have made. You should not only look at the textual information but also consider the input visual information, taking everything into account. Base on the provided information, your task is to explain the reason behind your choice.
The frame tokens: {Frame tokens}
The question is: {Question}
The options is:{Options}
The Key Info is:{Key Info}
Your choice is {Choice}
Prompt for confidence Evaluation You are given a single-choice question, options, some frames of the long video, your choice for the question and the reason. Your task is to evaluate whether the selected choice is correct. Criteria for Evaluation: Insufficient Information (Confidence Level: 1): There isn’t enough clear information to be sure. Partial Information (Confidence Level: 2): The decision is likely correct, but there’s some uncertainty. Sufficient Information (Confidence Level: 3): The decision is clearly supported by the given information. You need to provide the Confidence Level (a number of 1 or 2 or 3) according to the Criteria for Evaluation. Choose the confidence level that best reflects how clearly the input supports your choice. Respond with only the JSON format: ’confidence’: ’x’, where ’x’ is Confidence Level, without any additional text, explanation, or formatting.
The frame tokens: {Frame tokens}
The question is: {Question}
The options is:{Options}
Your choice is {Choice}
Your reason is {Reason}
Prompt for updating the key information in Round i Given some frames from a long video, subtitles, a single-choice question, and options, the old key information and the old choices and reasons, identify the key information needed to answer the question. Focus on visual cues, context, and temporal relationships within the frames. Limit your response to 50 words.
The frame tokens: {Frame tokens}
The question is: {Question}
Round 1 Key Info: {Key Info in Round 1}
Round 1 Choice: {Choice in Round 1}
Round 1 Reason: {Reason in Round 1}
Round i-1 Key Info: {Key Info in Round i-1}
Round i-1 Choice: {Choice in Round i-1}
Round i-1 Reason: {Reason in Round i-1}
Appendix B Visualization and Experimental Comparison of Different Models.
As shown in Fig. 5, direct reasoning models such as GPT-4o [34] and Qwen2.5VL-72B [1] perform global sampling on videos and struggle to focus on key information in long videos. Agent-based methods like VideoTree [46] often suffer from incorrect or redundant sampling to generate noisy caption for wrong answer, due to the lack of deep thinking on shots in a long video. Although VideoAgent [43] implements a dynamic frame selection process, it performs retrieval in a frame-by-frame manner, which makes it susceptible to redundant information. Moreover, VideoAgent relies heavily on captions to answer questions, resulting in the loss of substantial visual information. While CoS [13] shares the same name—Chain of Shot—with the method proposed in this work, it does not implement a multi-round dynamic shot partition. Instead, it instead performs only a single-step inference. Specifically, it uniformly samples frames from the video and groups every four adjacent frames into a single “shot.” All shots are then categorized into positive and negative shots based on their relevance to the given question. The model generates three separate answers: one from the positive shots, one from the negative shots, and one from uniformly sampled frames. These responses are aggregated to produce the final answer. In contrast, VideoChat-A1 interactively employs shot perception and reasoning via Chain-of-Shot, which progressively looks into relevant shots through a reflective process to achieve superior performance.
Table 9 presents the experimental results of our method compared with the agent-based methods shown in Fig. 5. VideoChat-A1 (InternVL2.5-8B) achieves 72.1% on Egoschema, corresponding to absolute gains of 18.0% and 11.0% over VideoAgent (54.1%) and VideoTree (61.1%), respectively. On LongVideoBench, VideoChat-A1 (InternVideo2.5-8B) attains 65.4%, improving by 12.6% over CoS (LongVA) and by 6.5% over CoS (LLaVA-Video). For MLVU (M-avg), VideoChat-A1 (InternVideo2.5-8B) reaches 76.2%, outperforming CoS (LongVA) by 17.3% and CoS (LLaVA-Video) by 4.8%. Finally, on Video-MME without subtitles, VideoChat-A1 (InternVideo2.5-8B) achieves 72.9%, with 19.4% and 8.5% improvements over CoS (LongVA) and CoS (LLaVA-Video). On Video-MME with subtitles, it achieves 76.8%, with 21.1% and 5.3% improvements over CoS (LongVA) and CoS (LLaVA-Video), respectively.
Model Egoschema LongVideoBench MLVU Video-MME
(M-avg) Short Medium Long Average
VideoAgent [43] 54.1 - - - - - -
VideoTree (GPT-4)  [46] 61.1 - - - - - -
CoS (LongVA[13] - 52.8 58.9 61.6/64.2 52.0/54.4 46.8/48.5 53.5/55.7
CoS (LLaVA-Video[13] - 58.9 71.4 77.2/80.1 62.4/69.4 53.8/65.1 64.4/71.5
Table 9: Comparison with other three agent-baesd models on four Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs".
Figure 5: Visual comparison of different models.
Appendix C Details of the Training of LongCLIP.
To enable LongCLIP to handle long-video understanding tasks, we construct a fine-tuning dataset by sampling videos from ActivityNet [17], OpenVid-1M [33], ViTT [15], and MovieChat-Caption [40]. Specifically, we first select videos that are already segmented into multiple clips, each paired with corresponding captions. These clip-caption pairs are directly used as training samples. For the remaining videos, we filter out those shorter than 5 seconds, longer than 15 minutes, or whose captions contain fewer than 10 words. After filtering, we manually create a subset, resulting in a long-video retrieval dataset consisting of 47k samples. The average video length is 101.40 seconds, and each caption contains an average of 72.29 words. The composition of the dataset is summarized in Table 10.
Dataset Instance Num Avg Caption Length Avg Video Length
ActivityNet-Caption 2,1352 15.20 165.32
OpenVid-1M 2,2348 132.22 8.78
ViTT 3,024 22.51 280.04
MovieChat-Caption 496 132.92 432.57
Total 47,220 72.29 101.40
Table 10: Instance numbers of different datasets for training the LongCLIP.
Since LongCLIP is originally designed for image inputs, we adapt it to video by uniformly sampling 16 frames from each shot and applying a 3D convolution to extract spatiotemporal features compatible with LongCLIP’s input format. We adopt the training hyperparameters following the LongCLIP paper, with a batch size of 1024, 3 training epochs, a weight decay of 1e-2, and a learning rate of 1e-4.

### v2

@@ Introduction
We evaluate VideoChat-A1 on four long-video question-answering benchmarks: EgoSchema (Mangalam et al. 2023), LongVideoBench (Wu et al. 2024), MLVU (Zhou et al. 2024), and Video-MME (Fu et al. 2024). As baselines, we use Qwen2.5-VL-7B, InternVL2.5-8B, and InternVideo2.5-8B. In the Feature Extraction stage, we sample the video at 1 fps and use pretrained CLIP-ViT-B/32 (Radford et al. 2021) as the feature extractor. In Shot Selection, we sample 16 frames from each shot and use a fine-tuned LongCLIP-B (Zhang et al. 2024a) to calculate shot-text similarity. The details are described in the supplementary documentation. We retrieve shots with similarity scores above the 0.8 threshold. In the first round, we treat the whole video as the candidate shot without selection, and sample $\mathcal{T}=16$ frames as initial frames. In each subsequent round, we select top 2 shots, divide each shot into 2 subshots, and sample $\mathcal{T}=8$ frames from each subshot, and add them into frame set. In the Shot Partition stage, the number of clusters $K$ is set to 6 in 1st round and 2 in subsequent rounds. The maximum number of iterations is 3. All experiments are conducted on 2 A800-80GB GPUs.
Figure 4: Shot Reasoning and Shot Reflection. At each round, VideoChat-A1 performs question-answering reasoning using the relevant shots identified. It then evaluates the confidence level of the response. Then, the system either proceeds to the next step for further refinement or terminates the reasoning process to output the final answer.
Comparison with SOTA
Performance and Efficiency Analysis.
@@
As shown in Table 1, we evaluate VideoChat-A1 on four mainstream benchmarks. Specifically, VideoChat-A1 significantly outperforms open-source 7-8B models, comparable to or exceeding that of large open-source or closed-source models. For example, on MLVU, VideoChat-A1 (InternVideo2.5-8B) achieves 76.2%, outperforming GPT-4o by 11.6%. Table 2 compares the average frame number and inference time. VideoChat-A1 achieves substantial gains in efficiency. For instance, VideoChat-A1 (InternVL2.5-8B) reduces frame usage by over 90% and inference time by more than 85% compared to GPT-4o, Gemini 1.5 Pro, and Qwen2.5-VL-72B. Despite these significant savings, our method maintains competitive performance across benchmarks. In contrast to small open-source models, VideoChat-A1 not only delivers significantly better performance but also requires fewer video frames and achieves faster or comparable inference.
Ablation Study
@@
Table 3: Ablation on the Video Glance in VideoChat-A1.
@@
In this study, we conducted an ablation study to verify the impact of the video glance. Table 3 shows that the model with the video glance achieves higher scores on all the tested tasks. The improvements proving the effectiveness of video glance.
@@
In Table 4, we evaluate the impact of Shot Partition (SP) and Chain-of-Shot (CoS) on VideoChat-A1 (Qwen2.5-VL-7B). The EgoSchema and LongVideoBench are abbreviated as Ego and LVbench in the table because of the limited space. In the setup without SP, we evenly divided the entire video into six shots on average in the 1st round. In subsequent rounds, for the selected shots that need to be expanded, they are directly partitioned into two shots on average. Regarding the setup without CoS, we removed the CoS mechanism. Instead, the model samples additional frames from the video and incorporates them into the existing frame set from the previous round. Then we conducted a vote to get the final response. Individually, enabling SP or CoS boosts performance. Their combined use yields the best results. For example, on MLVU, the score rises from 66.2 (with neither) to 71.9 (both enabled).
SP CoS Ego LVBench MLVU VideoMME
@@
Table 4: Ablation of Shot Partition (SP) and Chain-of-Shot (CoS) in VideoChat-A1 (Qwen2.5-VL-7B).
@@
Table 5: Ablation of different methods of shot partition in VideoChat-A1 (Qwen2.5-VL-7B).
@@
Table 7: Ablation over number of initialized shots in VideoChat-A1 (Qwen2.5-VL-7B).
@@
As shown in Table 5, we evaluate the impact of three methods of shot partition across multiple benchmarks. Cluster Partition refers to the clustering approach used in VideoTree (Wang et al. 2024c), where all frames within each cluster are treated as a single shot, even though they may not form a continuous video segment. Average Partition refers to dividing each shot into smaller shots by uniformly splitting it into equal parts. Our proposed Shot Partition method consistently outperforms the other two methods, which demonstrate that partitioning videos into temporally coherent shots based on semantic information leads to better understanding of long videos.
@@
We evaluate the impact of varying the maximum number of reflection rounds. As shown in Table 6, increasing the number of rounds consistently leads to performance improvements. In the context of multi-round discussions, increasing the number of discussion rounds tends to enhance the effectiveness. After three rounds of the chain-of-shot process, the results are already quite satisfactory. As the fourth round yields only marginal improvements, we ultimately adopt the three-round setting to balance performance gains and computational cost.
@@
We evaluate the effect of the number of initial shots, which is set to 3, 6, and 9. As shown in Table 7, the 6-shot achieves best accuracies compared with the other two configurations. For the subsequent rounds, considering that iterative process generates many subshots and aiming to balance computational efficiency, we fix clustering number for each subshot at $K=2$.
Conclusion
@@
In this paper, we introduced VideoChat-A1, an agent-based framework for effectively addressing long video understanding tasks. Unlike existing methods that often overlook the intrinsic shot-based structure , our method explicitly employs a Chain-of-Shot reasoning paradigm. Specifically, VideoChat-A1 progressively identifies and refines relevant video segments through three key iterative steps: Shot Selection, Shot Partition, and Shot Reflection. This iterative interaction mechanism closely mimics human cognitive processes, allowing the agent to thoughtfully reason and accurately answer the questions. Experimental results demonstrate the effectiveness and efficiency of our method, highlighting the substantial benefits in long video understanding tasks.
Acknowledgements
Supported by Shanghai Artificial Intelligence Laboratory, and the National Key R&D Program of China(NO.2022ZD0160505).